# DET-LSH: A Locality-Sensitive Hashing Scheme with Dynamic Encoding Tree for Approximate Nearest Neighbor Search

### v1

@@ 5.3. Complexity Analysis
In the query phase, DET-LSH has time cost $\mathcal{O}(n(\beta d+\log N_{r}))$. The time cost comes from four parts: (1) computing hash values for the query point $q$, $\mathcal{O}(L\cdot K\cdot d)=\mathcal{O}(d)$; (2) finding candidate points in $L$ DE-Trees, $\mathcal{O}(L\cdot K^{2}\cdot\log N_{r}\cdot\frac{n}{max\_size}+L\cdot K\cdot n)=\mathcal{O}(n\log N_{r})$; (3) computing the real distance of each candidate point to $q$, $\mathcal{O}(\beta nd)$; and (4) finding the $top$-$k$ points to $q$, $\mathcal{O}(\beta n\log k)$. The total time cost in the query phase is $\mathcal{O}(n(\beta d+\log N_{r}))$.
@@ 6.2.2. Query Phase
In practice, in Algorithm 5, if the upper bound distance between a leaf node and $q^{\prime}$ is greater than the search radius, it will take much time to calculate the distance between each point in the leaf node and $q^{\prime}$ (lines 8-13). After experiments, we found that if the leaf node size $max\_size$ is appropriately set in Algorithm 3, most of the points in these “troublesome” leaf nodes are within the search radius. Therefore, we optimized Algorithm 5 in two aspects: (1) We relax the requirements for candidate points to improve efficiency. In our implementation, as long as the lower bound distance between a leaf node and $q^{\prime}$ is not greater than $r$, we will add all its points to $S$. (2) We maintain a priority queue to hold traversed leaf nodes based on their lower bound distances to $q^{\prime}$. A leaf node with a smaller lower bound distance to $q^{\prime}$ can add all its points to $S$ earlier, guaranteeing the quality of candidate points. As shown in Figure 5, with an acceptable sacrifice of query accuracy, optimized Algorithm 4 and Algorithm 7 improve query efficiency by up to 50% and 30%.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under the grant number 62202450.

### v2

@@ 5.3. Complexity Analysis
In the query phase, DET-LSH has time cost $\mathcal{O}(n(\beta d+\log N_{r}))$. The time cost comes from four parts: (1) computing hash values for the query point $q$, $\mathcal{O}(L\cdot K\cdot d)=\mathcal{O}(d)$; (2) finding candidate points in $L$ DE-Trees, $\mathcal{O}(L\cdot K^{2}\cdot\log N_{r}\cdot\frac{n}{max\_size}+L\cdot K\cdot n)=\mathcal{O}(n\log N_{r})$; (3) computing the real distance of each candidate point to $q$, $\mathcal{O}(\beta nd)$; and (4) finding the toptop-$k$ points to $q$, $\mathcal{O}(\beta n\log k)$. The total time cost in the query phase is $\mathcal{O}(n(\beta d+\log N_{r}))$.
@@ 6.2.2. Query Phase
In practice, in Algorithm 5, if the upper bound distance between a leaf node and $q^{\prime}$ is greater than the search radius, it will take much time to calculate the distance between each point in the leaf node and $q^{\prime}$ (lines 8-13). After experiments, we found that if the leaf node size max_sizemax\_size is appropriately set in Algorithm 3, most of the points in these “troublesome” leaf nodes are within the search radius. Therefore, we optimized Algorithm 5 in two aspects: (1) We relax the requirements for candidate points to improve efficiency. In our implementation, as long as the lower bound distance between a leaf node and $q^{\prime}$ is not greater than $r$, we will add all its points to $S$. (2) We maintain a priority queue to hold traversed leaf nodes based on their lower bound distances to $q^{\prime}$. A leaf node with a smaller lower bound distance to $q^{\prime}$ can add all its points to $S$ earlier, guaranteeing the quality of candidate points. As shown in Figure 5, with an acceptable sacrifice of query accuracy, optimized Algorithm 4 and Algorithm 7 improve query efficiency by up to 50% and 30%.
@@ Acknowledgements.
This work is supported by the National Natural Science Foundation of China (NSFC) under the grant number 62202450, and supported by the Hellenic Foundation for Research and Innovation (HFRI) under the “Second Call for HFRI Research Projects to support Faculty Members and Researchers” (project number: 3684).